1. An electrically neutral glass rod is rubbed against a neutral piece of silk. During this process 2.0 x 1015 electrons are transferred from the rod to the silk.
a. How much charge (in coulombs) did the silk acquire?
b. What charge does the glass rod have after being rubbed on the silk? (explain)
1
Expert's answer
2020-04-20T10:15:32-0400

a). Silk has acquired a charge

"Q=e \cdot N"

Where "N=2 \cdot 10^{15}" -the number of electrons that were transferred from the rod to silk.

"e=-1.6 \cdot10^{-19}C" -charge of one electron.

Then

"Q=e \cdot N=-1.6 \cdot10^{-19} \cdot 2 \cdot 10^{15}=-3.2 \cdot 10^{-4}C=-0.32 mC"

b). Prior to the friction of the glass rod on silk, the number of positive and negative charges was equally divided.

With the transfer of 0 electrons to silk, the number of positive charges on the rod became greater. That is, the rod acquired a charge "Q=+0.32 mC"